OpenStudy (anonymous):

Circumference of the circle A is 24 times pie and circumference of the circle B is12 times pie . What is the difference between the areas of A and B

OpenStudy (anonymous):

C = 2pi * r radius of circle A = 24pi = 2pi * r r = 12 radius of circle B = 12pi = 2pi * r r = 6 A = pi * r^2 Area of circle A = pi * 12^2 Area of circle B = pi * 6^2 A - B = 144pi - 36pi = 108pi

OpenStudy (anonymous):

thnx =)